OpenStudy (anonymous):

What is the solution to the equation 9^(x - 2) = 27

OpenStudy (lgbasallote):

hint: 9 = 3^2 27 = 3^3 does that help?

OpenStudy (anonymous):

um kinda

OpenStudy (anonymous):

x = 2.5 x = 3.5 x = 0.5 x = -0.5 are my answer choices

OpenStudy (anonymous):

Can someone help me?

OpenStudy (lgbasallote):

so apply my hint to the problem..

OpenStudy (anonymous):

I don't understand? I thought it would be 5 but thats not a answer choice

OpenStudy (anonymous):

or 27 = 9^(3/2)

OpenStudy (anonymous):

9^3 = 729

OpenStudy (lgbasallote):

okay so you have \[\Large 9^ {x-2} = 27\] apply my hint here (3^2 = 9; 3^3 = 27) \[\Large \implies (3^2)^{x-2} = 3^3\] does that help?

OpenStudy (anonymous):

so 3x^-4 = 9?

OpenStudy (anonymous):

i mean 3x^4 = 28

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

3 times x = 3x and 2 times -2 is negative 4 and 3^3 is 27

OpenStudy (lgbasallote):

why did you multiply 3 by x?

OpenStudy (anonymous):

bevause x is on the outside of the parenthesis

OpenStudy (lgbasallote):

hmm well you did something close to what was supposed to happen you apply the rule \[\large (a^m)^n = a^{mn}\] so that means \[\Large (3^2)^{x-2} = 3^3\] becomes\[\Large \implies 3^{2x - 4} = 3^3\] got it?

OpenStudy (anonymous):

kinda

OpenStudy (lgbasallote):

so now use the rule \[\Large a^x = a^y \Longleftrightarrow x = y\]

OpenStudy (anonymous):

so the answer is 3.5

OpenStudy (lgbasallote):

sounds right

OpenStudy (anonymous):

thanks for the help

OpenStudy (lgbasallote):

welcome ^_^